The Energy of Electrons in Heavy Atoms According to Jansen and Hess

Heinz Siedentop Ludwig-Maximilians-Universität München

We show self-adjointness for a relativistic one-electron Hamiltonian used to describe electrons in heavy atoms. The model is a approximate no-pair Hamiltonian in the Furry picture. (Joint work with Edgardo Stockmeyer and Raymond Brummelhuis.)